How many mLs of H2 gas, collected by water displacement, are produced by the reaction of 0.500 g of Mg metal with an excess of HCl, at an atmospheric pressure of 750 torr and a temperature of 25.0 C? The vapor pressure of water at 25.0 C is 23.8 torr.
1
Expert's answer
2013-05-16T11:06:43-0400
Thechemical equation of this process is
2Mg + 2HCl = MCl2 + H2
n(Mg) = 0.5 g / 24 g/moles = 0.021moles
n(H2) = 0.021moles *1/2 = 0.0105moles
pressure of dry H2 gas = 750 torr - 23.8torr = 726.2 torr
from the equation
PV = nRT
V = nRT/P
V(H2) = 0.0105moles * 8.314 * 298 K / 96 820 Pa = 0.27 mLAnswer: 0.27 mL